There are 45 out of the 54 people in writer's club who are left handed. What percent of the people are right handed?
1
Expert's answer
2020-12-10T10:54:05-0500
"\begin{matrix}
 54 & 100\% \\
 45 & x\%
\end{matrix}"

"x\%=\dfrac{45}{54}\cdot100\%=\dfrac{250}{3}\%"

"100\%-\dfrac{250}{3}\%=\dfrac{50}{3}\%\approx16.6667\%"

There are approximately 16.6667% right handed people among 54 people in writer's club.